OpenStudy (anonymous):

I reeaaaally need help, whats and equation for, (2,-3), y= -2x -3, that passes through the given point and is (a) parallel to the given line and (b) perpendicular to the given line.

OpenStudy (bahrom7893):

GO FOR IT MASTER SATELLITE! I don't even know what title level 100 is lol

OpenStudy (anonymous):

slope of \[y=-2x-3\] is \[-2\] right? so your job is to find the equation for the line through (2,-3) that has slope -2 do you know the point-slope formula??

OpenStudy (bahrom7893):

LEGEND Sattellite

OpenStudy (anonymous):

mistake maker satellite is more like it

OpenStudy (bahrom7893):

*satellite

OpenStudy (bahrom7893):

satellite, i need help with lin algebra. I'll take care of this, can u take a look at this question. The asker left, i'm tryin to get someone to help

OpenStudy (anonymous):

point - slope \[y-y_1=m(x-x_1)\] where \[m=-2,x_1=2,y_1=-3\]

OpenStudy (bahrom7893):

just finish ur work then i guess

OpenStudy (anonymous):

plug the numbers into the formula and you will get \[y-(-3)=-2(x-2)\] \[y+3=-2x+4\] \[y=-2x+1\]